The self-employment tax for a sole proprietor who reports $154,800 of self-employment profits on Schedule C and has no wages is:

$4,146

$21,873

$23,684

$20,068

$7,420

$7,763
Answer #1

Answer is option C

c. $20,068

Net earnings from self-employment = $154,800 x 92.35% = $142,958

Self-employment tax = (128400*12.4%)+(142958*2.9%) = 15922+4146 = $20068